OpenStudy (anonymous):

A spherical weather balloon is being inflated. The radius of the balloon is increasing at the rate of 7 cm/s. Express the surface area of the balloon as a function of time t (in seconds). (Let S(0) = 0.) S(t) = cm?

OpenStudy (campbell_st):

Use \[dSA/dt = dSA/dr \times dr/dt\] dr/dt =7 \[SA = 4 \pi r^2\] \[d(SA0=) = 8 \pi r\] \[d(SA)?dt = 8 \pi r \times 7\]

OpenStudy (campbell_st):

then \[d(SA)/dt = 56 \pi r\]

OpenStudy (anonymous):

nope

OpenStudy (campbell_st):

if dr/dt = 7 then by integrating r = 7t + c \[SA = 4 \pi r^2\] SA = \[SA = 4 \pi ( 7t + c)^2\]